# Symplectic Neural Flows for Modeling and Discovery

### v1

@@ 2.3 Training the SympFlow
$\mathcal{L}(\bar{\psi})=\mathcal{L}_{1}(\bar{\psi})+\mathcal{L}_{2}(\bar{\psi}),$ (12)
with $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ defined as in Equation 9 and Equation 10. For the MLP, we replace $\mathcal{L}_{2}$ with ~$\widetilde{\mathcal{L}}_{2}$. Notice that $\mathcal{H}(\bar{\psi})$ is the exact Hamiltonian behind the SympFlow $\bar{\psi}$, and the loss $\mathcal{L}_{2}(\bar{\psi})$ provides a comparison between the exact flow $\phi_{H,t}$ and the SympFlow via zeroth-order information, i.e., based on the exact energies behind those maps. Importantly, the $\mathcal{L}_{2}$ term allows us to perform a backward error analysis—while SympFlow does not generally solve the ODE under consideration, it solves a time-dependent Hamiltonian ODE, for a Hamiltonian which is driven towards $H$ during the training process. The fact that $\mathcal{H}(\bar{\psi})$ is time-dependent implies that the flow $\bar{\psi}$ does not conserve $\mathcal{H}$, as it is expected for non-autonomous Hamiltonian systems. However, if (10) is small enough, the Hamiltonian $\mathcal{H}(\bar{\psi})$ will not be strongly dependent on time, hence leading to an almost conservation of $H$ by the SympFlow. We further expand on this aspect below.
@@ 3 Theoretical analysis of the SympFlow
This section provides a theoretical a-posteriori analysis of a SympFlow approximating the target map $\phi_{H,t}$. This result relies on the assumption that the map $\phi_{H,t}$ can be accurately approximated for $t\in[0,\Delta t]$ by a SympFlow. This is possible since SympFlows are universal approximators in the space of Hamiltonian flows. We state this result in 1, and prove it in Appendix D.
@@
Let $H:\mathbb{R}\times\Omega\to\mathbb{R}$, $\Omega\subset\mathbb{R}^{2d}$ compact, be a continuously differentiable function. For any $\varepsilon>0$, there is a SympFlow $\bar{\psi}_{t}$ such that
@@
The details can be found in Appendix D. ∎
Remark 1.
For the a-posteriori analysis, we assume that there is a compact subset $\Omega\subset\mathbb{R}^{2d}$ which is forward invariant both for the flow $\phi_{H,t}$ we are approximating and for the network $\bar{\psi}_{t}$. More explicitly, we assume that $\bar{\psi}_{t}(x_{0})\in\Omega$ for every $x_{0}\in\Omega$ and $t\in[0,\Delta t]$, and $\phi_{H,t}(x_{0})\in\Omega$ for every $x_{0}\in\Omega$ and $t\geq 0$.
Based on the time extension provided in Equation 7 and Remark 1, one can obtain a function which accurately approximates a target flow map $\phi_{H,t}:\Omega\to\Omega$ for every time $t\geq 0$. This time extension is again the composition of Hamiltonian flows, hence possessing an underlying time-dependent Hamiltonian. For a fixed time, $t\geq 0$, the function in Equation 7 satisfies
@@
almost everywhere. We provide details about the derivation of Equation 13 in Appendix B. From Equation 13, we see that the long-time extension of a SympFlow is again the solution of a non-autonomous Hamiltonian system with time-dependent Hamiltonian function $\mathcal{H}(\psi):[0,+\infty)\times\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ defined as
@@
Therefore, the Hamiltonian in Equation 14 is piecewise continuous, and this is because $\psi$ is not differentiable at the time instants $t_{k}=k\Delta t$, $k\in\mathbb{N}$.
@@
Let $\bar{\psi}:[0,\Delta t]\times\Omega\to\Omega$ be a SympFlow. Let us assume that for every $x\in\Omega$ and $t\in[0,\Delta t]$
@@
$\left|\mathcal{H}({\psi})(0,x)-\mathcal{H}(\psi)(t,\psi_{t}(x))\right|\leq c_{1}\varepsilon_{1}t,$ (17)
$\left|H(x)-H(\psi_{t}(x))\right|\leq c_{2}(\varepsilon_{1}+\varepsilon_{2}t),$ (18)
where $\psi$ is defined as in Equation 7, and $\mathcal{H}(\psi)$ as in Equation 14.
We prove 2 in Appendix C. We remark that Equation 18 can be obtained with similar techniques for any neural network satisfying Equation 16, while Equation 17 only in the case one has a SympFlow and some a-posteriori bound as in Equation 15. Furthermore, we also point out that the assumptions in Equation 15 and Equation 16 correspond to saying that the loss function in Equation 12 is smaller than a specific constant, hence why this is an a-posteriori error estimate.
@@ 4 Numerical Experiments
The proposed experimental analysis compares the results obtained with SympFlow and an unconstrained neural network, which we will refer to as MLP. We provide some details on the network architectures in Appendix E.
We consider three test problems: the simple harmonic oscillator, the damped harmonic oscillator, and the Hénon–Heiles system. Details on these Hamiltonian systems are provided in Section 4.1, Section 4.2, and Section 4.3. For each test problem, we first consider the unsupervised task of solving the equations of motion. We then move on to consider the supervised task of estimating the flow based on irregular trajectory samples. The training process for these experiments follows the steps presented in Section 2.3.
@@
As we will see across all the simulations, SympFlow leads to considerably improved long-time energy behavior compared to MLP.
@@ 4.1.1 Unsupervised experiments
In this section, we compare the SympFlow architecture with an unconstrained MLP on solving the equations of motion in Equation 20. We train the SympFlow in three different ways, aiming to identify the impact of the two terms in the loss function, see Equation 12. The MLP is instead always trained with the regularization term in Equation 11. In the figures, the term regularization refers to the addition of the Hamiltonian matching term in Equation 10. When a SympFlow is trained with Hamiltonian regularization, it means it is trained to minimize the loss in Equation 12. If there is no such regularization, it means that we only train with the residual-based loss function in Equation 9. We also consider a mixed training procedure, which consists of a first training phase with Hamiltonian regularization, and a later fine-tuning of the weights by training the model for a few more epochs without Hamiltonian matching. Figure 3 reports the results obtained with these four training configurations. We see that the SympFlow architecture consistently outperforms the unconstrained MLP in predicting the correct qualitative behavior of the solution. This is evident both looking at the reproduced orbit, which are all associated to the initial condition $(q_{0},p_{0})=(1,0)$, but also in the long-time energy behavior. We remark that the orbit of $(1,0)$ is wholly contained in $\Omega$, so the networks are expected to approximate it reliably. The second column in Figure 3 shows the variation of the true Hamiltonian energy in Equation 19 along the network-produced solutions. Even though we extrapolate over the long integration time $[0,1000]$, the SympFlow outperforms the MLP regardless of how it has been trained. This is a consequence of the symplectic nature of the network. In this experiment, it is unclear whether the three different considered training procedures for the SympFlow have a significant impact. We explore this further in the other two test problems.
@@ 4.1.2 Supervised experiments
We now move to the supervised experiments, where we recall that the models are trained based on the mean squared error loss function in Equation 8. We first show in Figure 4 the obtained results with $N=100$ initial conditions, each sampled at possibly different $M=50$ time instants. In this case, we assume there is no noise, i.e., $\varepsilon=0$. The results in Figure 4 show a similar pattern to the unsupervised experiments, where the SympFlow outperforms the MLP over long time simulations. Comparing this figure with Figure 3, we notice that the unsupervised case leads to a smaller energy variation than the supervised one for SympFlow predictions. This difference is expected since the unsupervised experiment, even though it does not rely on trajectory data, relies on the knowledge of the Hamiltonian system one is trying to solve, hence having access to the true Hamiltonian function while training the network.
@@
We now evaluate the impact of the three parameters $N$, $M$, and $\varepsilon$ on the performance of SympFlow and MLP. We denote with $\psi:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ the extension of the network $\bar{\psi}$ over the real line, defined as in Equation 7. To quantitatively evaluate the results and compare the models, we consider the average relative error
$\frac{1}{I}\sum_{i=1}^{I}\frac{\left\|\psi(k\Delta t,x_{i}^{0})-\phi_{H,k\Delta t}(x_{i}^{0})\right\|_{2}}{\left\|\phi_{H,k\Delta t}(x_{i}^{0})\right\|_{2}},\,\,k=1,10,100$
over $I=100$ randomly sampled initial conditions in $\Omega$. We remark that even though the quantity above is expressed using the exact flow map $\phi_{H,t}$, in practice, we replace it with a reference numerical solution obtained using a Runge–Kutta $(5,4)$ integrator with tight tolerance. We also test the average relative Hamiltonian energy variation
$\frac{1}{I}\sum_{i=1}^{I}\frac{\left|H(\psi(k\Delta t,x_{i}^{0}))-H(x_{i}^{0})\right|}{\left|H(x_{i}^{0})\right|},\,\,k=1,10,100.$
We collect the values of these two metrics in Figure 5. Each of the three subfigures lets one of the three parameters vary and fixes the other two. We now discuss some outcomes from this experimental analysis:
@@
  • The relative error at $100\Delta t$ for the two models is comparable in all of the configurations. This is expected for two reasons:
@@
      The fact that SympFlow is symplectic does not ensure improved quantitative long-time behavior, but qualitative. This is also why one would not expect a symplectic time integrator to be more accurate than a classical one. However, as for the energy behavior above, the qualitative properties of the produced solutions are considerably improved.
  • Even in the presence of noise, the SympFlow outperforms MLP. This can be seen by looking at the error at time $\Delta t$ for different values of $\varepsilon$ in Figure 5(c). The symplectic constraint on SympFlow prevents it from overfitting the data and learning the noise.
  • The MLP looks to benefit from additional data much more than SympFlow. This can be seen in the relative errors and energy variations as $N$ and $M$ increase. This can be interpreted as a consequence of the fact that MLP has no encoded structure, and to infer the correct behavior, it needs a considerable amount of data.
(a) Relative $\ell^{2}-$norms of the error and the energy variations at times $\Delta t,10\Delta t,100\Delta t$ for a fixed $N=100$ and $\varepsilon=0$.
(b) Relative $\ell^{2}-$norms of the error and the energy variations at times $\Delta t,10\Delta t,100\Delta t$ for a fixed $M=50$ and $\varepsilon=0$.
(c) Relative $\ell^{2}-$norms of the error and the energy variations at times $\Delta t,10\Delta t,100\Delta t$ for a fixed $N=100$ and $M=50$.
Figure 5: Supervised experiment — Simple Harmonic Oscillator: Impact of the parameters $N$, $M$, and $\varepsilon$ on the training procedure.
@@ 4.2 Modelling Dissipation: Damped Harmonic Oscillator
Real-world dynamical systems often dissipate energy to their surroundings through mechanisms such as heat, friction, or radiative losses. These processes reduce the energy of the system until it reaches equilibrium. To accurately capture the dynamics of dissipative physical systems, it is critical to account for those effects within their equations of motion. To apply our approach to non-conservative systems, we adopt the formulation introduced in [23], which allows to express their dynamics as conservative systems in a phase space having doubled dimension compared to the physical space. This formalism enables to accommodate dissipation while retaining key features of the Hamiltonian dynamics, such as symplecticity, and hence apply SympFlow in a meaningful way. For more details on this formalism, see Appendix A. In this formulation, the degrees of freedom are doubled, and a point in this augmented space is given by $z_{a,b}:=(q_{a},q_{b},\pi_{a},\pi_{b})\in\mathbb{R}^{4d}$, where $\pi_{a}$ and $\pi_{b}$ are the so-called non-conservative momenta. The Hamiltonian of the augmented system is given by the function
@@
where $H:\mathbb{R}^{2d}\to\mathbb{R}$ is the Hamiltonian of the system, and $K:\mathbb{R}^{4d}\to\mathbb{R}$ is an interaction term modeling the action of the non-conservative forces. In our experiments, the function $K$ takes the form [55, 24]
$K(q_{a},q_{b},\pi_{a},\pi_{b})=-\frac{\lambda}{2}(\dot{q}_{a}+\dot{q}_{b})(q_{a}-q_{b}),$
giving rise to a damped harmonic oscillator. Although a trajectory described by the original phase-space variables, $(q(t),p(t))$, does not follow a conservative dynamics, the trajectory described in the augmented space, is indeed conservative. Hence, in the augmented space, the map $(t,z_{a,b}(0))\mapsto z_{a,b}(t)$ is symplectic and the equations of motion can be recast in a symplectic form:

### v2

@@ 2.3 Training the SympFlow
$\mathcal{L}(\bar{\psi})=\mathcal{L}_{1}(\bar{\psi})+{\gamma}\mathcal{L}_{2}(\bar{\psi}),\,\,{\gamma\in\{0,1\}},$ (13)
with $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ defined as in Equation˜10 and Equation˜11. For the MLP, we replace $\mathcal{L}_{2}$ with $\widetilde{\mathcal{L}}_{2}$. In the experiments, we compare the case without regularization, $\gamma=0$, with the one with regularization, $\gamma=1$. Such a regularization significantly improves the results for the MLP. This does not occur for SympFlow, which performs well in the unregularized case as well, which, sometimes, also outperforms the regularized network. Notice that $\mathcal{H}(\bar{\psi})$ is the exact Hamiltonian behind the SympFlow $\bar{\psi}$, and the loss $\mathcal{L}_{2}(\bar{\psi})$ provides a comparison between the exact flow $\phi_{H,t}$ and the SympFlow via zeroth-order information, i.e., based on the exact energies behind those maps. Importantly, the $\mathcal{L}_{2}$ term allows us to perform a backward error analysis—while SympFlow does not generally solve the ODE under consideration, it solves a time-dependent Hamiltonian ODE, for a Hamiltonian which is driven towards $H$ during the training process. The fact that $\mathcal{H}(\bar{\psi})$ is time-dependent implies that the flow $\bar{\psi}$ does not conserve $\mathcal{H}$, as it is expected for non-autonomous Hamiltonian systems. However, if Equation˜11 is small enough, the Hamiltonian $\mathcal{H}(\bar{\psi})$ will not be strongly dependent on time, hence leading to an almost conservation of $H$ by the SympFlow. We further expand on this aspect below.
@@ 3 Theoretical analysis of the SympFlow
This section provides a theoretical a-posteriori analysis of a SympFlow approximating the target map $\phi_{H,t}$. This result relies on the assumption that the map $\phi_{H,t}$ can be accurately approximated for $t\in[0,\Delta t]$ by a SympFlow. This is possible since SympFlows are universal approximators in the space of Hamiltonian flows. We state this result in ˜1, and prove it in Appendix˜F.
@@
Let $H:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}$ be a twice-continuously differentiable function, and fix $\Omega\subset\mathbb{R}^{2d}$ compact. Suppose that $\Omega$ is forward invariant for $\phi_{H,t}$. For any $\varepsilon>0$, there is a SympFlow $\bar{\psi}_{t}$ such that
@@
The details can be found in Appendix˜F. ∎
Based on the time extension provided in Equation˜8 and Remark˜1, one can obtain a function which accurately approximates a target flow map $\phi_{H,t}:\Omega\to\Omega$ for every time $t\geq 0$. This time extension is again the composition of Hamiltonian flows, hence possessing an underlying time-dependent Hamiltonian. For a fixed time, $t\geq 0$, the function in Equation˜8 satisfies
@@
almost everywhere. We provide details about the derivation of Equation˜14 in Appendix˜A. From Equation˜14, we see that the long-time extension of a SympFlow is again the solution of a non-autonomous Hamiltonian system with time-dependent Hamiltonian function $\mathcal{H}(\psi):[0,+\infty)\times\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ defined as
@@
Therefore, the Hamiltonian in Equation˜15 is piecewise continuous, and this is because $\psi$ is not differentiable at the time instants $t_{k}=k\Delta t$, $k\in\mathbb{N}$.
@@
Let $H:\mathbb{R}^{2d}\to\mathbb{R}$ be a twice-continuously differentiable function, $\Delta t>0$, $\Omega\subset\mathbb{R}^{2d}$ compact, and $\bar{\psi}:[0,\Delta t]\times\Omega\to\Omega$ be a SympFlow. Suppose that $\Omega$ is forward invariant for both $\phi_{H,t}$ and $\bar{\psi}_{t}$, i.e., for every $x_{0}\in\Omega$ and $t\in[0,\Delta t]$ one has $\phi_{H,t}(x_{0}),\bar{\psi}_{t}(x_{0})\in\Omega$. Assume that for every $x\in\Omega$ and $t\in[0,\Delta t]$
@@
$\left|\mathcal{H}({\psi})(0,x)-\mathcal{H}(\psi)(t,\psi_{t}(x))\right|$ $\leq{c_{1}(\varepsilon_{1}+\varepsilon_{2}t)},$ (18)
$\left|H(x)-H(\psi_{t}(x))\right|$ $\leq{c_{2}\varepsilon_{2}t},$ (19)
where $\psi$ is defined as in Equation˜8, and $\mathcal{H}(\psi)$ as in Equation˜15.
Remark 1.
Forcing $\Omega$ to be forward invariant for $\bar{\psi}_{t}$ before training is extremely challenging. Still, since this is an a-posteriori analysis, it assumes that the network has already been trained to match the exact flow, and almost-forward invariance is what we observe experimentally. Such a property is a consequence of a successful training procedure coupled with the symplectic nature of SympFlow. Indeed, its symplecticity experimentally leads to almost energy conservation and, consequently, to almost forward invariance of $\Omega$.
We prove ˜2 in Appendix˜B. We remark that Equation˜19 can be obtained with similar techniques for any neural network satisfying Equation˜17, while Equation˜18 only in the case one has a SympFlow and some a-posteriori bound as in Equation˜16. Furthermore, we also point out that the assumptions in Equation˜16 and Equation˜17 correspond to saying that the loss function in Equation˜13 is smaller than a specific constant, hence why this is an a-posteriori error estimate.
@@ 4 Numerical Experiments
The proposed experimental analysis compares the results obtained with SympFlow and an unconstrained neural network, which we will refer to as MLP. We provide some details on the network architectures in Appendix˜C.
We consider three test problems: the simple harmonic oscillator, the damped harmonic oscillator, and the Hénon–Heiles system. Details on these Hamiltonian systems are provided in Section˜4.1, Section˜4.2, and Section˜4.3. For each test problem, we first consider the unsupervised task of solving the equations of motion. We then move on to consider the supervised task of estimating the flow based on irregular trajectory samples. The training process for these experiments follows the steps presented in Section˜2.3. A PyTorch implementation of our architecture and these experiments can be found at the repository https://github.com/davidemurari/sympflow.
@@
As we will see across all the simulations, SympFlow leads to considerably improved long-time energy behavior compared to MLP. Furthermore, the Hamiltonian regularization considerably benefits the MLP for the more challenging problems such as Hénon-Heiles, but does not improve the already good performance of the unregularized SympFlow.
We provide a thorough evaluation of SympFlow with additional experiments in the appendix. In Appendix˜I, we compare the computational time of the training and inference phases for SympFlow with that of the MLP. We test networks with a varying number of layers $L\in\{4,8,16\}$, report the loss curves, and also compare the quality of the obtained approximation. Our experiments in this section of the main body of the paper focus on a comparison of SympFlow with an MLP, which is the strategy proposed in [43] to approximate $\phi_{H,t}$. We also compare our SympFlow with the Symplectic Recurrent Neural Network procedure presented in [14]. These experiments are collected in Appendix˜J and showcase SympFlow’s robustness to noise.
@@ 4.1.1 Unsupervised experiments
In this section, we compare the SympFlow architecture with an unconstrained MLP on solving the equations of motion in Equation˜21. We train the SympFlow in three different ways, aiming to identify the impact of the two terms in the loss function, see Equation˜13. The MLP is instead either trained with or without the regularization term in Equation˜12, i.e., with a residual term alone or coupled with Hamiltonian regularization. When a network is trained with Hamiltonian regularization, it means it is trained to minimize the loss in Equation˜13 with $\gamma=1$. If there is no such regularization, it means that we only train with the residual-based loss function in Equation˜10, i.e., $\gamma=0$ in Equation˜13. In Appendix˜G, we also consider a mixed training procedure consisting of a first training phase with Hamiltonian regularization, and a later fine-tuning of the weights by training the model for a few more epochs without Hamiltonian matching. Figure˜3 reports the results obtained with these four training configurations. We see that the SympFlow architecture consistently outperforms the unconstrained MLP in predicting the correct qualitative behavior of the solution. This is evident both looking at the reproduced orbit, which are all associated to the initial condition $(q_{0},p_{0})=(1,0)$, but also in the long-time energy behavior. We remark that the orbit of $(1,0)$ is wholly contained in $\Omega$, so the networks are expected to approximate it reliably. The second column in Figure˜3 shows the variation of the true Hamiltonian energy in Equation˜20 along the network-produced solutions. Even though we extrapolate over the long integration time $[0,1000]$, the SympFlow outperforms the MLP regardless of how it has been trained. This is a consequence of the symplectic nature of the network. In this experiment, it is unclear whether the different training procedures considered have a significant impact. We explore this further in the other two test problems.
@@ 4.1.2 Supervised experiments
We now move to the supervised experiments, where we recall that the models are trained based on the mean squared error loss function in Equation˜9. We first show in Figure˜4 the obtained results with $N=100$ initial conditions, each sampled at possibly different $M=50$ time instants. In this case, we assume there is no noise, i.e., $\varepsilon=0$. The results in Figure˜4 show a similar pattern to the unsupervised experiments, where the SympFlow outperforms the MLP over long time simulations. Comparing this figure with Figure˜3, we notice that the unsupervised case leads to a smaller energy variation than the supervised one for SympFlow predictions. This difference is expected since the unsupervised experiment, even though it does not rely on trajectory data, relies on the knowledge of the Hamiltonian system one is trying to solve, hence having access to the true Hamiltonian function while training the network.
@@
We now evaluate the impact of the three parameters $N$, $M$, and $\varepsilon$ on the performance of SympFlow and MLP. We denote with $\psi:\mathbb{R}\times\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ the extension of the network $\bar{\psi}$ over the real line, defined as in Equation˜8. To quantitatively evaluate the results and compare the models, we consider the average relative error
$\frac{1}{I}\sum_{i=1}^{I}\frac{\left\|\psi(k\Delta t,x_{i}^{0})-\phi_{H,k\Delta t}(x_{i}^{0})\right\|_{2}}{\left\|\phi_{H,k\Delta t}(x_{i}^{0})\right\|_{2}},\,\,k=1,10,100$ (22)
over $I=100$ randomly sampled initial conditions in $\Omega$. We remark that even though the quantity above is expressed using the exact flow map $\phi_{H,t}$, in practice, we replace it with a reference numerical solution obtained using a Runge–Kutta $(5,4)$ integrator. We also test the average relative Hamiltonian energy variation
$\frac{1}{I}\sum_{i=1}^{I}\frac{\left|H(\psi(k\Delta t,x_{i}^{0}))-H(x_{i}^{0})\right|}{\left|H(x_{i}^{0})\right|},\,\,k=1,10,100.$ (23)
We collect the values of these two metrics in Figure˜5. Each of the three subfigures lets one of the three parameters vary and fixes the other two. We now discuss some outcomes from this experimental analysis:
@@
  • The relative error for the two models is comparable in many configurations, with SympFlow generally achieving lower errors. This is expected for two reasons:
@@
      The fact that SympFlow is symplectic does not ensure improved quantitative (long-time) behavior, but qualitative. This is also why one would not expect a symplectic time integrator to be more accurate than a classical one. However, as for the energy behavior above, the qualitative properties of the produced solutions are considerably improved.
  • Even in the presence of noise, the SympFlow outperforms MLP. This can be seen by looking at the error at time $\Delta t$ for different values of $\varepsilon$ in Figure˜5(c). The symplectic constraint on SympFlow prevents it from overfitting the data and learning the noise.
  • Both models benefit from additional data as $N$ and $M$ increase, with the MLP showing a stronger dependence on data in several settings. This can be interpreted as a consequence of the fact that the MLP has no encoded structure, and to infer the correct behavior, it needs a considerable amount of data.
(a) Fixed $N=100$ and $\varepsilon=0$.
(b) Fixed $M=50$ and $\varepsilon=0$.
(c) Fixed $N=100$ and $M=50$.
Figure 5: Supervised experiment — Simple Harmonic Oscillator: Relative $\ell^{2}$-norms of the error and the energy variations at times $\Delta t,10\Delta t,100\Delta t$.
@@ 4.2 Modelling Dissipation: Damped Harmonic Oscillator
Real-world dynamical systems often dissipate energy to their surroundings through mechanisms such as heat, friction, or radiative losses. These processes reduce the energy of the system until it reaches equilibrium. To accurately capture the dynamics of dissipative physical systems, it is critical to account for those effects within their equations of motion. To apply our approach to non-conservative systems, we adopt the formulation introduced in [25], which allows to express their dynamics as conservative systems in a phase space having doubled dimension compared to the physical space. This formalism enables to accommodate dissipation while retaining key features of the Hamiltonian dynamics, such as symplecticity, and hence apply SympFlow in a meaningful way. For more details on this formalism, see Appendix˜E. In this formulation, the degrees of freedom are doubled, and a point in this augmented space is given by $z_{a,b}:=(q_{a},q_{b},\pi_{a},\pi_{b})\in\mathbb{R}^{4d}$, where $\pi_{a}$ and $\pi_{b}$ are the so-called non-conservative momenta. The Hamiltonian of the augmented system is given by the function
@@
where $H:\mathbb{R}^{2d}\to\mathbb{R}$ is the Hamiltonian of the system, and $K:\mathbb{R}^{4d}\to\mathbb{R}$ is an interaction term modeling the action of the non-conservative forces. In our experiments, the function $K$ takes the form [63, 24] $K(q_{a},q_{b},\pi_{a},\pi_{b})=-\lambda/2(\dot{q}_{a}+\dot{q}_{b})(q_{a}-q_{b})$, giving rise to a damped harmonic oscillator. Although a trajectory described by the original phase-space variables, $(q(t),p(t))$, does not follow a conservative dynamics, the trajectory described in the augmented space, is indeed conservative. Hence, in the augmented space, the map $(t,z_{a,b}(0))\mapsto z_{a,b}(t)$ is symplectic and the equations of motion can be recast in a symplectic form: